# Unifying Explainable Anomaly Detection and Root Cause Analysis in Dynamical Systems

### v1

@@ Introduction
For our theoretical analysis, we assume that for a period of time, $Z$ does not change. Unlike measurement anomalies, cyber anomalies directly alter variables within the ODE, modifying the system’s evolution and potentially propagating to dependent variables. Such anomalies can be caused by cyber-attacks, including false data injection (Sayghe et al. 2020), which fundamentally disrupt the system’s dynamics.
@@
$C_{i,j}=median_{0\leq t\leq T-1}\left\{|\Phi_{\theta}(X(t))|\right\},\textit{for }1\leq i,j\leq p.$ (9)
@@
The calculation process involves calculating the proportion of largest $m$ elements in the difference matrix $|C-C^{\prime}|$ for each row, and taking the maximum proportion as the measurement score $M(C,C^{\prime})$. A greater $M(C,C^{\prime})$ indicates a higher likelihood that the anomaly is a measurement anomaly, since the changes are concentrated on one column. While a lower score suggests a cyber anomaly. This classification is crucial for subsequent root cause localization, which we will discuss in the following section.
@@
Consider a dynamical system $X(t)$ as given in Eq. (1), where $\mathcal{G}=(X(t),\mathcal{E})$ denotes the underlying dependency graph. Let $C(\cdot,\cdot)$ and $C^{\prime}(\cdot,\cdot)$ denote the causality matrices extracted using non-anomalous and anomalous data, respectively. Assuming $C(i,j)\neq 0$ iff $(i,j)\in\mathcal{E}$: (1) If the anomaly is a measurement anomaly at variable $k$, then $C(i,j)-C’(i,j)\neq 0$ only if $k=i$. (2) If the anomaly is a cyber anomaly at variable $k$, then $C(i,j)-C’(i,j)\neq 0$ only if $i=k$ and $(i,j)\in\mathcal{E}$.
@@
These results support our hypothesis that learning the ODE in dynamical systems can improve anomaly detection and root cause localization. ICODE’s approach aligns with the mechanism of anomaly occurrence: exceptions due to ODE changes are located by learning about the ODE and inferring the root cause through changes in ICODE parameters. In contrast, baseline models lack this grounding in dynamical systems science. CausalRCA’s causal inference model may struggle to learn causal relationships in this context, while RootClam’s use of distribution changes for root cause localization may be less effective in dynamical systems. In the following part, we will discuss the anomaly type classification based on proposed theorems.

### v3

@@ Introduction
For our theoretical analysis, we assume that for a period of time, $A(\cdot)$ does not change. Unlike measurement anomalies, cyber anomalies directly alter variables within the ODE, modifying the system’s evolution and potentially propagating to dependent variables. Such anomalies can be caused by cyber-attacks, including false data injection (Sayghe et al. 2020), which fundamentally disrupt the system’s dynamics.
@@
$C_{i,j}=median_{0\leq t\leq T-1}\left\{|\Phi_{\theta,(i,j)}(X(t))|\right\},\textit{for }1\leq i,j\leq p.$ (9)
@@
The calculation process involves calculating the proportion of the largest $m$ elements in the difference matrix $|C-C^{\prime}|$ for each row, and taking the maximum proportion as the measurement score $M(C,C^{\prime})$. A greater $M(C,C^{\prime})$ indicates a higher likelihood that the anomaly is a measurement anomaly, since the changes are concentrated on one column. While a lower score suggests a cyber anomaly. This classification is crucial for subsequent root cause localization, which we will discuss in the following section.
@@
Consider a dynamical system $X(t)$ as learned using Eq. (5), where $\mathcal{G}=(X(t),\mathcal{E})$ denotes the underlying dependency graph. Let $C(\cdot,\cdot)$ and $C^{\prime}(\cdot,\cdot)$ denote the causality matrices extracted using non-anomalous and anomalous data, respectively. Assuming $C(i,j)\neq 0$ iff $(i,j)\in\mathcal{E}$, and the diagonal elements of $C$ are much larger than the off-diagonal elements, $C(i,i)\gg C(i,j),i\neq j$, then (1) If the anomaly is a measurement anomaly at variable $k$, then $|C(k,j)-C’(k,j)|\gg|C(k^{\prime},j)-C’(k^{\prime},j)|>0$, where $(k,k^{\prime})\in\mathcal{E}$. (2) If the anomaly is a cyber anomaly at variable $k$, then $|C(k,j)-C’(k,j)|>0$ only if $(k,j)\in\mathcal{E}$.
@@
These results support our hypothesis that learning the ODE in dynamical systems can improve anomaly detection and root cause localization. ICODE’s approach aligns with the mechanism of anomaly occurrence: exceptions due to ODE changes are located by learning about the ODE and inferring the root cause through changes in ICODE parameters. In contrast, baseline models lack this grounding in dynamical systems science. CausalRCA’s causal inference model may struggle to learn causal relationships in this context, while RootClam’s use of distribution changes for root cause localization may be less effective in dynamical systems. In the following part, we will discuss the anomaly type classification based on the proposed theorems.
@@ Conclusion
Acknowledgements
This work was sponsored by a Research Futures grant from Lehigh University and the Office of Naval Research (ONR) under grant number N00014-22-1-2626.